A particle is projected vertically upwards from a roof of a tower 60 m high with velocity 20 m/s. Find the average speed upto an instant when it strikes the ground. g = 10 m/s(sq)

Grade:11

2 Answers

vikas askiitian expert
13 years ago

let T1 be the time taken by the particle to reach at the highest point 

 applying v=u-gT1  , v=0at the highest point ,u=20m/sec

   T1=2sec 

   let S1 be the distance covered during this period 

  then S1 = uT1-(gT1^2)/2

        S1=20m  

        same time is taken by particle to come at the point of projection

    now the velocity of particle becomes 20m/sec in downward dyrection 

      let time taken by particle to raech the ground is T2,again applying eq of motion

     -60=-20T2-(gT2^2)/2

    T2=2sec 

   total time =2T1+T2=6sec

   total distance=60+2.20=100 

   average speed =100/6=16.67m/s

Akash
4 years ago
A block moves ina straight line with velocity v for time t°.then it's velocity becomes 2vfor next t° of ime finally it's velocity becomes 3v for timeT.if average velocity during the complete journey was2.5v then find T in terms of t°